OpenStudy (anonymous):

Last two questions for the day please help will medal, and as always thank you loves. solve for x log5x + log4 = 3 solve for x log12(x + 44) = 2

OpenStudy (anonymous):

first of all, do you know how to get rid of logs?

Nnesha (nnesha):

first one is totally same like last one if you understand last question then its going to be easy for u :)

Nnesha (nnesha):

addition change to ??? go to previous post and look d information that i gave u

OpenStudy (mathstudent55):

First, use this on the left side of the first equation. \(\log a + \log b = \log (ab) \)